OpenStudy (anonymous):

find the slope of the line connecting the following pairs of points 4. (14,9) & (14,0)

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

(0-9)/(14-14) =-9/0 slope DNE

OpenStudy (anonymous):

\[m = \frac{y{2} - y{1}}{x{2} - x{1}}\] plug in the numbers and you get \[m = \frac{0 - 9}{14 - 14}\] so you get \[m = -\frac{9}{0}\] soo its negative infinity lol wtf did i do something wrong??

OpenStudy (anonymous):

ok Agent got the same answer lol.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no plot the points man. its a strait line up and down

OpenStudy (anonymous):

Yeah^^^ there is no slope in a vertical line

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Dam tricky questions lol.. Medal for you.

OpenStudy (anonymous):